# Wealth Thermalization Hypothesis

### v2

@@ Abstract
Figure 2: Color plot of wealth $w$ from Lorenz curves of the RJS model ($N=10000$). The $x$-axis corresponds to the fraction of households $h\in[0,1]$ and the $y$-axis to the rescaled energy $\varepsilon=E/B\in[0,0.5]$. The ticks mark integer multiples of 0.1 for $h$ and $\varepsilon$.
RJ thermalization and condensation. We start from a model with $N$ equidistant energy levels $0\leq E_{m}=m/N\leq B$ located in the energy band of total width $B$. This corresponds to certain levels of wealth for $N$ agents with a fraction of agents on level $m$ being $\rho_{m}$. The conserved average system energy is $E=\sum_{m}E_{m}\rho_{m}$ and the dimensionless parameter $\varepsilon=E/B$ determines its fraction with respect to the maximal system energy $B$. We call this model the RJ standard (RJS) model. On the basis of WTH with RJ distribution (1) the local (normalized) wealth on level $i$ is $(E_{i}/E)\rho_{i}$ and the cumulated wealth on levels $[0,m]$ is $w=\sum^{m}_{i=0}(E_{i}/E)\rho_{i}$ with the cumulated fraction of population or households $h=\sum^{m}_{i=0}\rho_{i}$. Computing both sums for all values of $m=0,1,\ldots$ provides the Lorenz curve $w(h)$. Since the Lorenz curve describes the normalized distribution of cumulated fractions of wealth $w\in[0,1]$ and households $h\in[0,1]$ we use the ratio $E_{i}/E$ (since $E=\sum_{i}E_{i}\rho_{i}$) in the definition of wealth ensuring that $w=1$ at $h=1$ for the total population. At given $\varepsilon$ the relation (1) and two integrals of energy and norm determine the physical parameters $T,\mu,S$. In our numerical studies we use $N=10000$ which practically corresponds to the continuous limit with results independent of $N$. The dependencies of $T$ and $\mu$ on $\varepsilon$ in the RJS model are shown in Supplementary Material (SupMat) Fig. S1. As discussed in rmtprl ; ourfiber for $\varepsilon>1/2$ the temperature $T$ becomes negative and at $\varepsilon$ close to unity there is a formation of an RJ condensate on highest energy levels with $E_{m}\rightarrow B$ (see SupMat Fig. S2). Many unusual properties of RJ thermalization have been discussed in rmtprl ; ourfiber but for convenience we provide some details in SupMat and Figs. S2 and S3 show the dependence of $\rho_{m}$ on $E_{m}/B$ for certain values of $\varepsilon$ with a clear formation of an RJ condensate at small $\varepsilon$ or $1-\varepsilon$. Even if the regime with negative temperatures has been realized in fiber experiments picozzi3 we consider that such a regime is not applicable to human society and hence we consider only the range with $20\leq\varepsilon\leq 1/$ 2.
Figure 3: Comparison of the Lorenz curves for US 2019 (black), UK 2012-2014 (blue), World 2021 (dashed green) with those of RJS model (red curves; $N=10000$); US and World curves are rather close. For the three referenced curves Gini coefficients are $G=0.852$, $0.626$, $G=0.842$ respectively and the rescaled energies $\varepsilon=E/B$ of RJS model are respectively fixed as $\varepsilon=0.07420$, $\varepsilon=0.1996$, $\varepsilon=0.07911$ so that the corresponding Gini coefficients match the referenced data.
Figure 4: Both panels compare the Lorenz curves for different data sets (black for US 2019, blue for UK 2012-2014 and green dashed for World 2021) with those of the RMT model (a) and the DL model (b). As in Fig. 3 the Gini coefficients $G$ of the reference curves are used to fix the rescaled energy $\varepsilon=E/B$ of the corresponding model such that the model curves (red) have the same $G$. For the RMT model (a) only two data sets are shown $\varepsilon=0.07996$ (US) and $\varepsilon=0.2027$ (UK). For the DL model (b) the parameter values are $a=16$ (US and World) and $a=3$ (UK). These values are fixed to have a best possible fit of the model data with those of the reference curves. The chosen values $\varepsilon=0.01434$ (US), $\varepsilon=0.1355$ (UK), $\varepsilon=0.01535$ (World) match the $G$ values of the reference data. In (b) the curves for US and World are rather close and a zoomed view is given in SupMat Fig. S4. In (a) the RMT Lorenz curves are shown for one realization of a random matrix, other realizations give practically the same curves.
The Lorenz curves for the RJS model at several $\varepsilon$ values are shown in Fig. 1. Due to RJ condensate there is a very high fraction of poor households $f_{p}$ (with $w\leq 0.02$) and a small fraction of rich ones $f_{r}$ (with $w\geq 0.75$) who owns a huge fraction of total wealth. Thus the RJS model naturally describes the big phase of poor households and the oligarchic phase of small fraction of households capturing the big fraction of total wealth. At maximal $\varepsilon=0.5$ with ($\mu\to-\infty$) all $\rho_{m}$ are equal and hence the Lorenz curve is $w=h^{2}$ with the limiting minimal Gini coefficient $G=1/3$ for the RJS model. The dependence of cumulated wealth $w$ on $h$ and $\varepsilon$ is shown in Fig. 2. It clearly shows the phase of poor households (blue), corresponding to the RJ condensate, and the oligarchic phase of very rich households (red). Thus we see that the RJ thermal distribution (1) describes the main qualitative features of wealth inequality of human society piketty2 .
From Figs. 12 we see that for the RJS model the WTH based on (1) captures main elements of wealth inequality but it is important to see if it can reproduce the real Lorenz curves for the whole world and specific countries. For this comparison of WTH theory we choose three cases with the Lorenz curves for: the whole world from piketty2 (integration of front page data gives cumulative $w,h$ values); USA 2019 case from usa2019 and UK 2012-2014 case from uk2014 . These real Lorenz curves are compared with those obtained from the RJS model (1) in Fig. 3. For the comparison values of $\varepsilon$ are fixed in such a way that Gini coefficient is the same for theory and real data curves. The comparison for UK case shows that there is a good agreement of real and theoretical Lorenz curves even if there is a certain difference for the range $0.9\leq h\leq 1$. The difference is more visible for USA case and the whole world (Lorenz curves are very similar for these two cases). In SupMat Fig. S4, we also show that the the RJS Lorenz curves have a satisfactory agreement with the Lorenz curves for France and Germany (data are obtained for the year 2010 from defr ).
In view of certain differences between real Lorenz curves and those obtained from RJS model (see Fig. 3) we also study the case of RJ distribution (1) with level energies $E_{m}$ taken from a random matrix of size $N=1000$ as it was discussed in rmtprl . For this RJ RMT model the density of states is $\nu=dm/dE_{m}=\frac{2N}{\pi}\sqrt{1-E^{2}}$ with typical eigenvalues in the interval $E_{m}\in[-1,1]$ and we shift all $E_{m}$ to $E_{m}-E_{0}$ to have nonnegative values $E_{m}\geq 0$ in (1). The comparison of Lorenz curves for US and UK cases with the results of the RJ RMT model is shown in Fig. 4a. The similarity between real and RMT model data is a bit less good then those in Fig. 3 for the RJS model. This shows that the density of states $\nu$ can affect the Lorenz curves. Indeed, we have $\nu=$ const. for the RJS model being different from the semi-circle law of RMT model.
To reproduce the real Lorentz curves from piketty2 ; usa2019 ; uk2014 in a better way we also analyze a double-linear (DL) model with energies $E_{m}=m/N$ for $m<N/2$ and $E_{m}=E_{N/2}+a(m-N/2)/N$ for $m\geq N/2$ at $N=10000$ with $a=16$ ($B=8.5$) for US and World data, and $a=3$ $(B=2)$ for UK data. In this type of model the density of states takes not one but two values being $\nu=1$ and $\nu=1/a$. The existence of two $\nu$ values can correspond to a society where high wealth energy $E_{m}$ values are only accessible to very rich people whose density is lower compared to common people. The comparison of real Lorenz curves with those of the DL model is shown in Fig. 4b (and its zoomed version SupMat Fig. S5) demonstrating a better proximity between real Lorenz curves and those from the DL model as compared to the results of the RJS model in Fig. 3. However, the DL model has two fit parameters $a,\varepsilon$ while the RJS model has only one $\varepsilon$.
We also remind that for the RJS model the minimal Gini value is $G=1/3$ that is reached at maximal physical value of $\varepsilon=1/2$. Thus to have $G<1/3$, we need to significantly modify the density of states $\nu$. Indeed, we can obtain a perfect complete wealth and energy equipartition with $w=h$ and $G=0$ for the case when all $E_{m}=E_{0}$ values are the same. In this case, the integrals of energy and norm give only one conservation law and all states have the same energy and same population. A small spectrum modification to $E_{m}=E_{0}+m/N$ with a constant energy offset $E_{0}$ leads to Lorenz curves being closer to the diagonal with small Gini values $G<1/3$ and a finite slope $w(h)\approx[E_{0}/(\varepsilon+E_{0})]\,h$ at small $h$ (see SupMat with additional discussion of this model and related Fig. S6). We call this model the equipartition (EQI) model.
We also show color Figures analogous to Fig. 2 for various models discussed above (see SupMat Fig. S7).
Figure 5: Gini coefficient versus rescaled energy $\varepsilon=(E-E_{0})/(E_{N-1}-E_{0})$ for the RJS model (red), RMT model (green), DL model (blue; only for the case $a=16$), and the EQI model (pink for the offset $E_{0}=0.1$ and cyan for $E_{0}=1$; same values of $E_{0}$ are used in SupMat Fig. S6). The thin black lines show the values of $G=0.852$, $G=0.842$ and $0.626$ for the data of US 2019, World 2021 and UK 2014. The intersection of these lines with the red and green curves correspond to $\varepsilon$ values used in Figs. 34.
Figure 6: Dependence of fraction of poor households $f_{p}=h(2\%)$ (owning 2% of wealth) and fraction of rich oligarchic households $f_{r}=1-h(25\%)$ (owning 75% of wealth) on the rescaled energy $\varepsilon=E/B$ for the RJS model.
The dependence of the Gini coefficient $G$ on $\varepsilon$ is given in Fig. 5 for the different models. In global the results show that an increase of $\epsilon$ leads to a reduction of $G$. Also in Fig. 6, we show the dependence of fractions of poor $f_{p}$ and rich oligarchic $f_{r}$ households on $\varepsilon$ for the RJS model. Thus at $\varepsilon=0.07$ we have $f_{p}=0.73$ and $f_{r}=0.097$ that is close to the real values $f_{p}=0.53$(US), $0.5$(World) and $f_{r}=0.09$(US), $0.1$ (World) while for UK we have $f_{p}=0.32$, $f_{r}=0.28$ corresponding to a higher $\varepsilon\approx 0.21$. Furthermore Fig. 6 shows that the fraction of poor households can be significantly reduced and the fraction of rich households can be increased by increasing parameter $\varepsilon$, thus diluting the oligarchic phase.
Finally, we mention that for the RJS model it is possible to work out analytic expressions (at $N\to\infty$; see SupMat section III) for the Lorenz curve and other quantities that accurately match the numerical data (see SupMat Fig. S8). These expressions depend on $\mu$ and at small $\varepsilon\leq 0.2$ (with $\mu\approx 0$), we have $w(h)\approx e^{(h-1)/\varepsilon}$ and $G\approx 1-2\varepsilon$, matching 3 values of $G$ in Fig. 1.
Discussion. In this work, we use the WTH approach (1 to describe the wealth distribution in a closed system that may be a country or the whole world. Our main argument is that in such a system interaction of agents is described by nonlinear equations with the conservation of two integrals of motion being total number of agents (norm or total probability analogous to number of system particles) and total wealth (analogous to total system energy). Under these conditions the wealth sharing is described by the universal RJ thermal distribution (1) as it is the case for various physical systems mayer ; landau ; zakharovbook ; picozzi1 ; picozzi2 ; babin ; chris ; picozzi3 ; rmtprl ; ourfiber . The striking feature of RJ thermalization (1) is that at low system energy (low $\varepsilon$) there is the physical phenomenon of RJ condensation when a high fraction of total probability is located at lowest energy states that corresponds to the high fraction of poor households with very low wealth and also other small fraction of oligarchic households that monopolize a big fraction of total wealth. Thus according to the WTH the phenomenon of huge wealth inequality in the world piketty1 ; piketty2 finds a natural thermodynamic explication. We show that the WTH theory gives a good description of the Lorenz curves of US, UK and the whole world.
On the basis of WTH theory we see that a reduction of wealth inequality can be realized by an increase of rescaled system energy ($\varepsilon=e/B$). The simplest way to reach this is to reduce the global dispersion of wealth (given by $B$) that can be realized by a high taxation of high wealth revenues.
Universality of WTH theory is also confirmed by showing that it well describes the Lorenz curves of capitalization of companies at S&P500 of New York Stock Exchange (NYSE), London SE, Hong Kong SE (see SupMat section IV and Figs. S9-S12 with data from sp500 ; london ; hk ).
Furthmore, another interesting extension of the RJS model with state density $\nu=dm/dE_{m}=N/(1+aE_{m})$ allows to obtain a significantly improved agreement of the theoretical curves with the diverse data sets (see SupMat section V and Figs. S13-S17) where the parameter $a$ is determined by a fit of a virtual reconstructed spectrum extracted from the real Lorenz curves (see SupMat section VI and Fig. S18).
Acknowledgments: This work has been partially supported through the grant NANOX $N^{o}$ ANR-17-EURE-0009 in the framework of the Programme Investissements d’Avenir (project MTDINA).
Supplementary Material for
Wealth Thermalization Hypothesis
by K. M. Frahm and D. L. Shepelyansky
Laboratoire de Physique Théorique, Université de Toulouse, CNRS, UPS, 31062 Toulouse, France
Submitted June 19, 2025 (updated June 30, 2025) with additional figures, explanations and analytical formula.
I I. General features of the thermalization in the RJS model
@@
where $\rho_{m}$ is the time averaged squared amplitude and occupation probability of each oscillator. If the non-linear terms are sufficiently strong or if there is some weak coupling to an external system (which respects both constraints (S.1)) one can assume that the system thermalizes. Applying the framework of the grand canonical ensemble one introduces two parameters: temperature $T$ and chemical potential $\mu$ to satisfy both constraints (S.1) in average and it can be shown (see e.g. Ref. [25]) that
$\rho_{m}=\frac{T}{E_{m}-\mu}\quad,\quad T=\frac{E-\mu}{N}$ (S.2)
@@
where the expression for the temperature $T$ is obtained from $\sum_{m}(E_{m}-\mu)\rho_{m}=(E-\mu)$ which follows directly from (S.1). The chemical potential is determined (using standard numerical techniques) by solving the implicit equation:
$1=\frac{E-\mu}{N}\sum_{m=0}^{N-1}\frac{1}{E_{m}-\mu}$ (S.3)
@@
which allows for one physical solution of $\mu$ outside the energy interval $[E_{\rm min},E_{\rm max}]$ with either $\mu<E_{\rm min}$ ($T>0$) or $\mu>E_{\rm max}$ ($T<0$) (depending on the value of $E$ we have either $T<0$ or $T>0$) such that $\rho_{m}>0$. The data presented in this work were obtained by this procedure for different model spectra and certain values of $N=10000$ (or $N=1000$ for the RMT model). Concerning the RJS model we have also considered the cases $N=100$, $N=1000$ and verified that the obtained Lorenz curves are very close (in graphical precision).
@@
Figure S1: The left (right) panel shows the (rescaled) temperature $NT$ (the chemical potential $\mu$) versus the rescaled energy $\varepsilon=E/B$ for the RJS model $E_{m}=m/N$, $N=10000$. The dashed black lines in the right panel correspond to the values of $E_{0}=0$ and $B\approx 1$ showing that either $\mu<E_{0}$ (for $T>0$) or $\mu>B$ (for $T<0$).
@@
As illustration Fig. S1 shows for the RJS model with $E_{m}=m/N$, $m=0,1,\ldots,N-1$, $N=10000$ both $T$ and $\mu$ as a function of $\varepsilon=E/B$ (here $B=E_{\rm max}-E_{\rm min}\approx 1$). Note that the left panel shows the rescaled temperature $NT$ since typical numerical values of $T$ are $\sim 1/N$ due to the finite value of $B$ in our particular model. (We note that the construction of the Lorenz curve is independent of a global scaling factor one could apply to the energy levels.) The figure illustrates that $-\mu\to 0$ ($\mu\to-\infty$) for $\varepsilon\to 0$ ($\varepsilon\to 1/2$).
@@
Figure S2: Color plot of the coarse grained thermalized occupation probabilities $\rho_{m}=T/(E_{m}-\mu)=(E-\mu)/[N(E_{m}-\mu)]$ for the RJS model. The $x$-axis corresponds to the fraction $E_{m}/B\in[0,1]$ (left to right) and the $y$-axis to the rescaled energy $\varepsilon$ (top to bottom for increasing values). The tics indicate integer multiples of 0.1 for both quantities. The color values shown in the color bar correspond to the value of $\rho_{m}$ averaged over intervals of size $1/20$ (for $E_{m}/B$ on the $x$-axis) and computed for 21 values $\varepsilon=i/20$, $i=0,1,\ldots,20$ (for the $y$-axis; the minimal value $\varepsilon=0$ has been slightly enhanced and the maximal value $\varepsilon=1$ has been slightly reduced to have a stable computation of the thermalized $\mu$-value). To increase visibility of small values a non-linear color bar scale has been chosen (e.g. green color corresponds to $1/16$).
@@
Figure S3: Dependence of the thermalized occupation probabilities $\rho_{m}=T/(E_{m}-\mu)=(E-\mu)/[N(E_{m}-\mu)]$ on $E_{m}/B$ for the RJS model $E_{m}=m/N$, $N=10000$ and the same values of $\varepsilon=E/B$ used in Figure 1 of the main part.
@@
Using (S.3) one can show that $-\mu\approx E/(N-1)\ll E$ for very small energies $0<E\ll 1/N$ and in this particular case we have $\rho_{0}\approx 1$ (strong condensation) and other $\rho_{m}\sim E/(NE_{m})\ll 1/N$ (for $m>0$). With increasing values of $E$ (or $\varepsilon$) the values of “$-\mu$” increase and more probability is shifted to the other $\rho_{m}$ values for $m>0$. At $\varepsilon\approx 1/2$ we have very large values of “$-\mu$” (and of $T$) such that all $\rho_{m}\approx 1/N$ are uniformly constant. Further increase of $\varepsilon$ enters the regime of negative temperatures (with $\mu>E_{\rm max}$) with possible condensation at the last oscillator with $\rho_{N-1}\gg 1/N$ (in this work we do not insist on the regime of $T<0$). These features are visible in both figures Figs. S2 and S3 showing $\rho_{m}$ versus $E_{m}/B$ for different values of $\varepsilon$ (as color plot or curves in log-scale). The effect of condensation for small $\varepsilon$ with a finite probability $\rho_{0}\gg 1/N$ is clearly visible in both figures and qualitatively one could even say that it extends even up to $\varepsilon=0.2$ with $\rho_{0}=0.002495$ still being larger than $1/N$. However, here also some other values of $\rho_{m}$ with small $m$ are significantly larger than $1/N$ (as can be seen in S3 for the first 5% of modes with $\rho_{m}\geq 3/N$). Also the coarse grained average value at the first 5% of modes at $\varepsilon=0.2$ is roughly 0.35 times the maximal coarse grained value at $\varepsilon\approx 0$ (according to Fig. S2). This effect corresponds to (modest) condensation on several modes or a given small mode interval.
@@
When constructing the Lorenz curve we have $w=0$ for $h<\rho_{0}$ and in the presence of (strong) condensation there is a finite interval of households with no wealth at all. Even for modest condensation over several modes the wealth value is initially very low. This can also be seen in Fig. 1 where $w(h\leq 0.1)\approx 0$ for $\varepsilon=0.2$ showing the effect of modest condensation.

### v3

@@ II.7 Wealth inequality and Lorenz curves
Figure II.9: Color plot of wealth $w$ from Lorenz curves for the netscience network with $\kappa=0.5$, $\beta=10$, $N=379$. The $x$-axis corresponds to the fraction of households $h\in[0,1]$ and the $y$-axis to the rescaled energy $\varepsilon=(E-E_{1})/(E_{N}-E_{1})\in[0,\varepsilon_{C}[$ where $\varepsilon_{c}=0.39622$ is the critical value at which the transition from $T>0$ to $T<0$ appears. The ticks mark integer multiples of 0.1 for $h$ and $\varepsilon$.
Fig. II.9 shows a color density plot for the Lorenz curves for a continuous distribution of $\varepsilon\in[0,\varepsilon_{c}[$ essentially confirming the observations of Fig. II.8 the length $h_{0}$ of the initial interval (with $w(h)=0$) behaves roughly as $h_{0}\approx\varepsilon/\varepsilon_{c}$.
Figure II.10: Lorenz curves for the politician network with $\kappa=0.5$, $\beta=10$, $N=5908$ for different values of the rescaled energy $\varepsilon=(E-E_{1})/(E_{N}-E_{1})$. The $x$-axis corresponds the cumulated fraction of households ($h$) and the $y$-axis to the cumulated fraction of wealth ($w$). The largest value $\varepsilon=0.28$ is slightly below the critical value $\varepsilon_{c}=0.28297$ at which the transition from $T>0$ to $T<0$ appears. The Gini coefficients $G$ for all curves are G=0.9328,0.8321,0.6642,0.4963,0.3284,0.1605,0.06647G=0.9328,0.8321,0.6642,0.4963,0.3284,0.1605,0.06647 (bottom to top).
Figure II.11: Color plot of wealth $w$ from Lorenz curves for the politician network with $\kappa=0.5$, $\beta=10$, $N=5908$. The $x$-axis corresponds to the fraction of households $h\in[0,1]$ and the $y$-axis to the rescaled energy $\varepsilon=(E-E_{1})/(E_{N}-E_{1})\in[0,\varepsilon_{C}[$ where $\varepsilon_{c}=0.28297$ is the critical value at which the transition from $T>0$ to $T<0$ appears. The ticks mark integer multiples of 0.1 for $h$ and $\varepsilon$.
Figs. II.10 and II.11 are as Figs. II.8 and II.9 respectively but for the politician network. Now the critical value for the transition from positive to negative $T$ is $\varepsilon_{c}=0.28297$. The structure of the curves and the color plot are rather similar as for the netscience network but the effect of straight lines for $h>h_{0}$ is even a bit stronger which is certainly related to the larger initial energy gaps of the politician network.
The results of this section show that the RJ thermalization and condensation in social networks presented in the previous sections leads to the formation of an enormous phase of very poor households and a small fraction of rich households that owns a significant fraction of total system wealth. The presence of a significant energy gap in the energy spectrum of two social networks studied here enhance the fraction of the poor phase comparing to the other spectrum models considered in Part I. The question of how typical such energy gaps are for social networks requires further studies. Another feature of the considered social networks is a stronglly peaked density of states approximately at the middle of the energy band. As a result the energies $E_{m}$ are very flat in this energy region. This is rather different from the RMT model or RJS model of Part I where the density of states is approximately constant in this energy range. We suppose that the origin of this difference is related to the fact that in the considered networks we have links between the members of the same society layer or class: scientists linked to scientists, politicians to politicians. Probably this is a general feature of internet connections where there is little if any distinction between classes of network members. In a real human society there is some kind of natural society statification: factory workers are mainly linked with workers, peasants with peasants, businessmen with businessmen, aristocrats with aristocrats. This feature is well present in a real human society with its society classes and wealth gradient between classes (of course with fluctuations and relatively weak links between classes). Thus it is possible that the society networks should be revised and updated to include the above feature of human society.
II.8 Overview of social networks results
The presented studies of dynamical thermalization in social networks show that chaos in these systems lead to the RJ thermalized distribution if nonlinear interactions are above a certain chaos border. The time scale for onset of this RJ distribution is determined by the strength of the nonlinearity. This time scale can be relatively long. On a first glance this seems rather surprising in view of a rather small number of links to hope and connect any pair of nodes, with the Erdös number $N_{E}\approx 4-5$ [30, 31, 65]. However, such link transitions are provided only by the linear part of system Hamiltonian while only the nonlinear interactions lead to transitions between eigenmodes (see (II.7)) with eventual thermalization. Our results show that the RJ thermalization process in the social networks has close similarities with those in the NLIRM model [17]. Thus we expect that the Lyapunov exponents $\lambda_{m}$ and the thermalization time scale $t_{RJ}$ have a similar to [17] dependence on nonlinearity $\beta$ and number of oscillators (nodes) where it was found that the typical Lyapunov exponent values are $\lambda\sim\beta^{1.5}/N^{1.9}$. We expect that $t_{RJ}\propto 1/\lambda$ but further studies are required to confirm these dependencies.
The emergence of RJ condensate leads to a formation of an enormous phase with high fraction of total norm located at low energy, or wealth, states. This leads to a massive fraction of poor households in the social networks as it is well seen in the figures of the Lorenz curves in Section II.7. At the same time a significant part of total wealth is captured by a small oligarchic group of rich households. Thus the obtained results for dynamical thermalization in social networks provide a confirmation of WTH origin and highlight the problem of wealth inequality in human society from a new view point.
Discussion and conclusion
In 1955 Fermi, Pasta, Ulam and Tsingu performed the first numerical simulations of a chain of nonlinear oscillators with the aim to find a dynamical thermalization and energy equipartition between the degrees of freedom. However, this model happened to have various specific features so that no tendency to equipartition was found in 1955 [2]. To extend these studies a generic model of coupled oscillators was proposed in [17] on the basis of nonlinear perturbation of Random Matrix Theory showing that chaos leads to dynamical thermalization with the resulting RJ distribution over the linear energy eigenmodes. This model has two integrals of motion being the total energy and total norm (probability). Thus the RJ distribution in this isolated Hamiltonian system is characterized by the system temperature $T(E)$ and chemical potential $\mu(E)$.
In fact the emergence of the RJ thermalization had been studied earlier numerically and experimentally for light propagation in multimode optical fibers [26, 48, 49, 50, 51, 52, 53], even if the origin of this thermalization was attributed to the Kolmogorov-Zakharov turbulence [25] without links to chaos and KAM integrability. The emergence of RJ condensation was established numerically [49] and experimentally [48, 50]. The emergence of an RJ condensate and thermalization in quantum chaos fibers with the nonlinear Schrödinger equation was demonstrated in numerical and analytical studies reported in [27]. At the same time it should be pointed out that the Fröhlich condensate for molecules at room temperature, discussed in [28, 29], has also certain similarities with the RJ condensate, even if in [28, 29] the system is considered under external pumping and dissipation (see discussion in [27]).
In Part I we analyzed the consequences of RJ thermalization and condensation associating system energy and norm, both conserved by time evolution. These quantities are related to the global wealth and the number of interacting households which are also conserved as justified in [45, 46]. The performed analysis shows that this WTH description depicts very well the shape of real Lorenz curves of wealth of households for several countries and the whole world. Also the WTH approach well reproduces the Lorenz curves for the stock exchange markets of New York, London and Hong Kong. To provide more arguments in support of the WTH description we study in Part II the dynamical thermalization in social networks induced by a nonlinear perturbation. Our results show the emergence of RJ thermalization and condensation in social networks with nonlinear interactions between network agents with the interactions being above a certain chaos border. As in Part I the RJ condensation leads to the Lorenz curves with an enormous fraction of poor households and a small fraction which owns a main part of total wealth. On the basis of the results of this work we argue that the WTH description provides new perspectives for the understanding of the nontrivial aspects of the wealth inequality in the world.
Declarations
  • Funding: The authors acknowledge support from the grant ANR France NANOX $N^{\circ}$ ANR-17-EURE-0009 in the framework of the Programme Investissements d’Avenir (project MTDINA).
  • Conflict of interest: The authors declare no conflicts of interest.
  • Data availability: This work uses specific datasets about wealth data for countries, stock market, network data etc. available at Refs. [31, 32, 33, 34, 37, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64].
[…]
Acknowledgements We thank A. D. Chepelianskii and L. Ermann for useful discussions. The authors acknowledge support from the grant ANR France NANOX $N^{\circ}$ ANR-17-EURE-0009 in the framework of the Programme Investissements d’Avenir (project MTDINA).
Appendix A Additional material for Part I
A.1 General features of the thermalization in the RJS model
@@
where $\rho_{m}$ is the time averaged squared amplitude and occupation probability of each oscillator. If the non-linear terms are sufficiently strong or if there is some weak coupling to an external system (which respects both constraints (A.1)) one can assume that the system thermalizes. Applying the framework of the grand canonical ensemble one introduces two parameters: temperature $T$ and chemical potential $\mu$ to satisfy both constraints (A.1) in average and it can be shown (see e.g. Ref. [17]) that
$\rho_{m}=\frac{T}{E_{m}-\mu}\quad,\quad T=\frac{E-\mu}{N}$ (A.2)
@@
where the expression for the temperature $T$ is obtained from $\sum_{m}(E_{m}-\mu)\rho_{m}=(E-\mu)$ which follows directly from (A.1). The chemical potential is determined (using standard numerical techniques) by solving the implicit equation:
$1=\frac{E-\mu}{N}\sum_{m=0}^{N-1}\frac{1}{E_{m}-\mu}$ (A.3)
@@
which allows for one physical solution of $\mu$ outside the energy interval $E_{\rm min},E_{\rm max}]$ with either $\mu<E_{\rm min}$ ($T>0$) or $\mu>E_{\rm max}$ ($T<0$) (depending on the value of $E$ we have either $T<0$ or $T>0$) such that $\rho_{m}>0$. The data presented in this work were obtained by this procedure for different model spectra and certain values of $N=10000$ (or $N=1000$ for the RMT model). Concerning the RJS model we have also considered the cases $N=100$, $N=1000$ and verified that the obtained Lorenz curves are very close (in graphical precision).
@@
Figure A.1: The left (right) panel shows the (rescaled) temperature NTNT (the chemical potential $\mu$) versus the rescaled energy $\varepsilon=E/B$ for the RJS model $E_{m}=m/N$, $N=10000$. The dashed black lines in the right panel correspond to the values of $E_{0}=0$ and $B\approx 1$ showing that either $\mu<E_{0}$ (for $T>0$) or $\mu>B$ (for $T<0$).
@@
As illustration Fig. A.1 shows for the RJS model with $E_{m}=m/N$, $m=0,1,\ldots,N-1$, $N=10000$ both $T$ and $\mu$ as a function of $\varepsilon=E/B$ (here $B=E_{\rm max}-E_{\rm min}\approx 1$). Note that the left panel shows the rescaled temperature $NT$ since typical numerical values of $T$ are $\sim 1/N$ due to the finite value of $B$ in our particular model. (We note that the construction of the Lorenz curve is independent of a global scaling factor one could apply to the energy levels.) The figure illustrates that $\mu\to 0$ ($\mu\to$-$\infty$) for $\varepsilon\to 0$ ($\varepsilon\to 1/2$).
@@
Figure A.2: Color plot of the coarse-grained thermalized occupation probabilities $\rho_{m}=T/(E_{m}-\mu)=(E-\mu)/[N(E_{m}-\mu)]$ for the RJS model. The $x$-axis corresponds to the fraction $E_{m}/B\in[0,1]$ (left to right) and the $y$-axis to the rescaled energy $\varepsilon$ (top to bottom for increasing values). The tics indicate integer multiples of 0.1 for both quantities. The color values shown in the color bar correspond to the value of $\rho_{m}$ averaged over intervals of size $1/20$ (for $E_{m}/B$ on the $x$-axis) and computed for 21 values $\varepsilon=i/20$, $i=0,1,\ldots,20$ (for the $y$-axis; the minimal value $\varepsilon=0$ has been slightly enhanced and the maximal value $\varepsilon=1$ has been slightly reduced to have a stable computation of the thermalized $\mu$-value). To increase visibility of small values a non-linear color bar scale has been chosen (e.g. green color corresponds to $1/16$).
@@
Figure A.3: Dependence of the thermalized occupation probabilities $\rho_{m}=T/(E_{m}-\mu)=(E-\mu)/[N(E_{m}-\mu)]$ on $E_{m}/B$ for the RJS model $E_{m}=m/N$, $N=10000$ and the same values of $\varepsilon=E/B$ used in Figure I.1 of the main part.
@@
Using (A.3) one can show that $\mu\approx E/(N-1)\ll E$ for very small energies $0<E\ll 1/N$ and in this particular case we have $\rho_{0}\approx 1$ (strong condensation) and other $\rho_{m}\sim E/(NE_{m})\ll 1/N$ (for $m>0$). With increasing values of $E$ (or $\varepsilon$) the values of $\mu$” increase and more probability is shifted to the other $\rho_{m}$ values for $m>0$. At $\varepsilon\approx 1/2$ we have very large values of $\mu$” (and of $T$) such that all $\rho_{m}\approx 1/N$ are uniformly constant. Further increase of $\varepsilon$ enters the regime of negative temperatures (with $\mu>E_{\rm max}$) with possible condensation at the last oscillator with $\rho_{N-1}\gg 1/N$ (in this work we do not insist on the regime of $T<0$). These features are visible in both figures Figs. A.2 and A.3 showing $\rho_{m}$ versus $E_{m}/B$ for different values of $\varepsilon$ (as color plot or curves in log-scale). The effect of condensation for small $\varepsilon$ with a finite probability $\rho_{0}\gg 1/N$ is clearly visible in both figures and qualitatively one could even say that it extends even up to $\varepsilon=0.2$ with $\rho_{0}=0.002495$ still being larger than $1/N$. However, here also some other values of $\rho_{m}$ with small $m$ are significantly larger than $1/N$ (as can be seen in A.3 for the first 5% of modes with $\rho_{m}\geq 3/N$). Also the coarse-grained average value at the first 5% of modes at $\varepsilon=0.2$ is roughly 0.35 times the maximal coarse-grained value at $\varepsilon\approx 0$ (according to Fig. A.2). This effect corresponds to (modest) condensation on several modes or a given small mode interval.
@@
When constructing the Lorenz curve we have $w=0$ for $h<\rho_{0}$ and in the presence of (strong) condensation there is a finite interval of households with no wealth at all. Even for modest condensation over several modes the wealth value is initially very low. This can also be seen in Fig. I.1 where $w(h\leq 0.1)\approx 0$ for $\varepsilon=0.2$ showing the effect of modest condensation.